OpenStudy (anonymous):

which ratio is equal to 4:9

geerky42 (geerky42):

8:18, 12:27, 16:36, etc

OpenStudy (anonymous):

just to let you know, in case you didn't already...ratios can be written as fractions. Example : 2:3 = 2/3 , 4:9 = 4/9 , 3:4 = 3/4 so basically your are just looking for an equivalent fraction to 4/9. you can find an equivalent fraction by multiplying numerator and denominator by the SAME NUMBER. 4/9 * 2/2 = 8/18 which can also be written like 8:18 4/9 * 3/3 = 12/27 which can be written like 12:27 Its that simple :)